# Towards the Worst-case Robustness of Large Language Models

### v3

@@ I.1 The Certified Bound is Still Weak
As analyzed in Sec. 5.1, the obtained $g(\bm{x}_{adv})$ for the absorbing kernel cannot exceed $\beta^{d}$. Since we typically set $\beta\leq 0.25$ and $d\geq 2$, it follows that $\beta^{d}\leq 0.1$. If we set the threshold $\tau\geq 0.1$, no theoretical guarantee can be obtained.
This limitation stems primarily from the formulation of Eq. 1. The current two knapsack solvers for Eq. 1 are indeed tight, i.e., there exists a worst-case bounded function $f$ for the fractional knapsack solver and a worst-case binary function $f$ for the 0-1 knapsack solver that satisfy all constraints in Eq. 1, with $g(\bm{x}_{adv})$ equal to the lower bound obtained by our solvers. In other words, the bound for Eq. 1 cannot be further improved. Since the worst-case model is excessively pessimistic, in the future, we may need to modify Eq. 1 to introduce additional constraints on the base model $f$ (e.g., Lipschitz continuity [52, 53]) to achieve a tighter bound.
In addition to revising the formulation of Eq. 1, certifying detectors rather than the base model itself offers an ad-hoc solution. For a detector, we can set the threshold $\tau$ as small as possible while ensuring a 0% false positive rate (FPR) on MTBench. Specifically, we choose $\tau=4.6\times 10^{-5}$ for $\beta=0.1$ and $\tau=4.6\times 10^{-4}$ for $\beta=0.25$. To validate the FPR on MTBench, we use a sample size of $N=100,000$ to estimate $g(\bm{x})=\mathbb{E}_{p(\bm{z}|\bm{x})}[f(\bm{z})]$. If the detector produces no false positives across these $N=100,000$ noisy samples, the confidence interval for the binomial proportion is $[0,4.6\times 10^{-5}]$. This justifies setting $\tau=4.6\times 10^{-5}$ for $\beta=0.1$.
@@ I.2 Upper Bound of Certified Radius due to Bayesian Error
$\boldsymbol{p_{A}}$ Bound. If the smoothing function $p(\bm{z}|\bm{x})$ removes all the keywords in $\bm{x}$, the subsequent model cannot produce the correct output. Thus, for uniform and absorbing kernel, the model accuracy is bounded as $p_{A}\leq 1-\beta^{K(\bm{x})}:=\overline{p_{A}}$. Consequently, we have:
@@ I.4 Defending against Expertise-based Attacks
The core principle of smoothing-based defenses is to transform out-of-distribution data back into in-distribution data, and its certified guarantees are effective only when the length of the adversarial suffix is limited. However, expertise-based attacks, which utilize human-crafted prompts, often appear natural (i.e., have high likelihood) and are typically lengthy, rendering our theoretical guarantees less effective (see ICA in Table 4). This issue could potentially be addressed by integrating our defense with existing heuristic defenses.
@@ I.5 Limited Settings of Certified Robustness
In this work, although we derive certifications for all smoothing distributions, there are still significant limitations. First, we cannot certify against heuristic attacks that use very long prompts, such as those in Wei et al. [36] and Chao et al. [11]. Additionally, we do not certify adversarial attacks involving insertion and deletion. This may require constructing $p(\bm{z}|\bm{x})$ to randomly insert or delete tokens. However, we believe that our framework can serve as a theoretical foundation, with future work focusing on proposing noising distributions of varying lengths and using fractional knapsack solver or 0-1 knapsack solver to certify against a broader class of attacks.
@@
Token consistency is simple in principle, but it took us a really long time to carefully ensure the token consistency for every model and defense, even each sentence. Of course, adaptive attacks are also crucial. One should at least include every part of the defense in the attacking process, rather than relying on techniques like BPDA [16]. Whether you design a specific loss function for each component, as in Carlini and Wagner [47], or treat the entire model as a unified procedure and optimize the overall loss does not make a significant difference.
@@
When applied to $\bm{x}_{\text{train}}$, we can certify that poisoning the training set may not significantly affect the functionality of the trained model, like Hong et al. [125]. When applied to $\bm{\theta}$, we can certify that corrupting or dropping out parts of $\bm{\theta}$ will not overly impact the functionality of the model or the training process. When applied to $\bm{x}_{\text{test}}$, as we have done, we can certify that adjusting the testing inputs will not successfully attack the already trained models.

### v4

@@ I.1 The Certified Bound is Still Weak
As analyzed in Sec.˜5.1, the obtained $g(\bm{x}_{adv})$ for the absorbing kernel cannot exceed $\beta^{d}$. Since we typically set $\beta\leq 0.25$ and $d\geq 2$, it follows that $\beta^{d}\leq 0.1$. If we set the threshold $\tau\geq 0.1$, no theoretical guarantee can be obtained.
This limitation stems primarily from the formulation of Eq.˜1. The current two knapsack solvers for Eq.˜1 are indeed tight, i.e., there exists a worst-case bounded function $f$ for the fractional knapsack solver and a worst-case binary function $f$ for the 0-1 knapsack solver that satisfy all constraints in Eq.˜1, with $g(\bm{x}_{adv})$ equal to the lower bound obtained by our solvers. In other words, the bound for Eq.˜1 cannot be further improved. Since the worst-case model is excessively pessimistic, in the future, we may need to modify Eq.˜1 to introduce additional constraints on the base model $f$ (e.g., Lipschitz continuity Chen et al. (2024a); Delattre et al. (2024)) to achieve a tighter bound.
In addition to revising the formulation of Eq.˜1, certifying detectors rather than the base model itself offers an ad-hoc solution. For a detector, we can set the threshold $\tau$ as small as possible while ensuring a 0% false positive rate (FPR) on MTBench. Specifically, we choose $\tau=4.6\times 10^{-5}$ for $\beta=0.1$ and $\tau=4.6\times 10^{-4}$ for $\beta=0.25$. To validate the FPR on MTBench, we use a sample size of $N=100,000$ to estimate $g(\bm{x})=\mathbb{E}_{p(\bm{z}|\bm{x})}[f(\bm{z})]$. If the detector produces no false positives across these $N=100,000$ noisy samples, the confidence interval for the binomial proportion is $0,4.6\times 10^{-5}]$. This justifies setting $\tau=4.6\times 10^{-5}$ for $\beta=0.1$.
@@ I.2 Upper Bound of Certified Radius due to Bayesian Error
$\bm{p_{A}}$ Bound. If the smoothing function $p(\bm{z}|\bm{x})$ removes all the keywords in $\bm{x}$, the subsequent model cannot produce the correct output. Thus, for uniform and absorbing kernel, the model accuracy is bounded as $p_{A}\leq 1-\beta^{K(\bm{x})}:=\overline{p_{A}}$. Consequently, we have:
@@ I.4 Defending against Expertise-based Attacks
The core principle of smoothing-based defenses is to transform out-of-distribution data back into in-distribution data, and its certified guarantees are effective only when the length of the adversarial suffix is limited. However, expertise-based attacks, which utilize human-crafted prompts, often appear natural (i.e., have high likelihood) and are typically lengthy, rendering our theoretical guarantees less effective (see ICA in Table˜4). This issue could potentially be addressed by integrating our defense with existing heuristic defenses.
@@ I.5 Limited Settings of Certified Robustness
In this work, although we derive certifications for all smoothing distributions, there are still significant limitations. First, we cannot certify against heuristic attacks that use very long prompts, such as those in Wei et al. (2023b) and Chao et al. (2023). Additionally, we do not certify adversarial attacks involving insertion and deletion. This may require constructing $p(\bm{z}|\bm{x})$ to randomly insert or delete tokens. However, we believe that our framework can serve as a theoretical foundation, with future work focusing on proposing noising distributions of varying lengths and using fractional knapsack solver or 0-1 knapsack solver to certify against a broader class of attacks.
@@
Token consistency is simple in principle, but it took us a really long time to carefully ensure the token consistency for every model and defense, even each sentence. Of course, adaptive attacks are also crucial. One should at least include every part of the defense in the attacking process, rather than relying on techniques like BPDA Athalye et al. (2018). Whether you design a specific loss function for each component, as in Carlini & Wagner (2017a), or treat the entire model as a unified procedure and optimize the overall loss does not make a significant difference.
@@
When applied to $\bm{x}_{\text{train}}$, we can certify that poisoning the training set may not significantly affect the functionality of the trained model, like Hong et al. (2024). When applied to $\bm{\theta}$, we can certify that corrupting or dropping out parts of $\bm{\theta}$ will not overly impact the functionality of the model or the training process. When applied to $\bm{x}_{\text{test}}$, as we have done, we can certify that adjusting the testing inputs will not successfully attack the already trained models.